OpenStudy (anonymous):

WILL FAN AND MEDAL :) Plz help me walk through this! Functions f(x) and g(x) are described as follows: f(x) = -2x^2 + 3 x g(x) 0 0 1 2 2 3 3 2 4 0

OpenStudy (anonymous):

Which statement best compares the maximum value of the two functions? It is equal for both functions. It is 2 units higher for f(x) than g(x). It is 2 units lower for f(x) than g(x). It is 1 unit higher for f(x) than g(x).

OpenStudy (anonymous):

@geerky42 @pooja195

OpenStudy (anonymous):

@Concentrationalizing @sammixboo

OpenStudy (anonymous):

@amistre64 @TheSmartOne

OpenStudy (anonymous):

@wio @IrishBoy123

OpenStudy (anonymous):

@Vocaloid @hartnn

OpenStudy (anonymous):

@Hero @perl

OpenStudy (anonymous):

@billj5 can you help me? :)

OpenStudy (anonymous):

Ok, so the part I'm having trouble with is finding the maximum of f(x)

OpenStudy (anonymous):

f(x) is a quadratic. The graph of a quadratic is a parabola, which would either have a finite minimum value or a finite maximum value. In this case you have a finite maximum value (because the question implies this and because the leading coefficient is negative). Given a quadratic with a general form of \(ax^{2} + bx + c = 0\), the x-coordinate of the maximum or minimum value is equal to \(-b/2a\).

OpenStudy (anonymous):

this might help http://www.purplemath.com/modules/fcntrans.htm

OpenStudy (anonymous):

Have you seen the -b/2a idea before, horse?

OpenStudy (anonymous):

Ok thank you @billj5 ;)

OpenStudy (anonymous):

Yes I have :) @Concentrationalizing

OpenStudy (anonymous):

Okay, awesome. So for our problem, what is -b/2a?

OpenStudy (anonymous):

I think it would be -3/2(-2)

OpenStudy (anonymous):

Actually, it would be 0. Note that the b in the general form is the coefficient of the x to the first power variable. Yet our equation has no x to the first power, we only have an x^2 and a constant term. Thus b is 0, which makes -b/2a = 0

OpenStudy (anonymous):

I meant those would be the values. So would the -3/2(-2) = 0?

OpenStudy (anonymous):

b is not -3, though, b is 0. Since a quadratic is \(ax^{2} + bx + c\), a is what goes with the x^2 term. So a for us is -2. b is what goes with the x term. But our equation is -2x^2 + 3, we dont have an x term. Thus b has to be 0. Then our constant, c, is 3. So a = -2, b = 0, c = 3. Which means -b/2a = 0/2(-2) = 0/-4 = 0 Does that make sense?

OpenStudy (anonymous):

But f(x) only has two terms?`:|

OpenStudy (anonymous):

So it's not a quadratic `:|

OpenStudy (anonymous):

A quadratic is any function in which all the powers are integers and the highest power is a 2. So all of these would be quadratics: \(x^{2} + 5x + 6\) \(\frac{-3x^{2}}{5} - 9\) \(x^{2}\) The first has 3 terms, the 2nd has 2, the last one has 1 term. But they're all quadratics because all of the exponents are integers and the highest one is a 2.

OpenStudy (anonymous):

Ok but I still don't get why b = 0.

OpenStudy (anonymous):